# Data Diversity Matters for Robust Instruction Tuning

### v1

@@ 4.2 Experimental Setup
QDIT:Threshold 10K 1.242 1.037 5.734 1.937 19.64 9.64 57.73
QDIT 10K 1.243 1.037 5.849 2.044 23.89 9.619 51.19
Winning Score vs. Alpaca 52K normal-$\uparrow$ Winning Score vs. Random 50K normal-$\uparrow$ HH Score Mean normal-$\uparrow$ Bottom 10% HH Score normal-$\uparrow$ Bottom 40% Lose Rate normal-$\downarrow$ Top 10% HH Scorenormal-$\uparrow$ Top 40% Win Rate normal-$\uparrow$
Random 50K 1.093 1 4.791 0.958 32.35 8.629 44.21
Quality 50K 1.22 1.182 5.963 2.133 30 9.361 53.06
QDIT 50K 1.217 1.168 6.074 2.278 34.48 9.484 51.55
Random 10K 1.108 1.002 4.906 0.981 26.21 8.67 44.66
Quality 10K 1.191 1.135 6.139 2.335 31.58 9.69 48.28
@@
QDIT:Threshold 10K 1.22 1.173 5.978 2.319 26.13 9.709 55.42
QDIT 10K 1.217 1.142 6.078 2.465 25.53 9.712 52.5
4.3 Main Results
The main experimental results can be found in Table 1. We mainly focus our analysis on the low data settings, as we find models trained with QDIT can match or beat the models trained in the high data settings while significantly reducing training cost. Nonetheless, we include the high data setting for reference.
Effect of Quality. Similar to prior works, we find that selecting data based on quality significantly improves average performance, improving the winning score versus Alpaca 52K by 5.41% and the average HH score by 17.14% in low data settings. However, we also find that selecting based on quality alone can hurt worst case performance, increasing the bottom 40% losing rate in two out of four settings compared to random data selection. We hypothesize that this performance drop is due to the fact that quality based selection hurts dataset diversity.
Effect of Diversity. On the other hand, we find that data selection with QDIT is able to achieve both a high average HH score (QDIT improves upon quality driven selection by 2.56% for winning score vs Alpaca 52K and improves average HH score by 1.29% in low data settings) while achieving a much better worst case performance than both random and quality-driven selection. In particular, we find that QDIT improves worst case HH score by 18.68 % and decreases worst case losing rate vs Alpaca 52K by 11.45% when compared to quality-driven data selection. We suspect that QDIT’s more diverse dataset teaches the model to a respond to a wide range of instructions, therefore decreasing the probability that it fails to follow evaluation instructions. We find that QDIT can also maintain best case performance when compared to quality-diven selection, as the best case HH score is improved by 0.12%. From these experiments we conclude that by increasing data diversity while maintaining data quality, QDIT can improve robust instruction following capability.
QDIT Variants. We also observe that the two variants of QDIT can improve average performance compared to random data selection, but they do not necessarily improve worst case performance compared to quality based selection. Our experiments suggest that clustering based data selection is highly sensitive to the performance of the clustering algorithm. For example, clustering may be easy on datasets like Alpaca which are derived from a small set of seed tasks, but difficult for larger and human generated instructions like LMSYS. For threshold based selection, we hypothesize that constraining intra-datapoint similarity is not as important as selecting instructions that cover the entire training set.
@@ 4.4 Analysis
Effect of $\alpha$. We plot the effect of $\alpha$ on average and worst case performance in Figure 3. From Figure 3, we can see that QDIT is relatively robust to the value of $\alpha$, with values of $\alpha\in\{0.5,0.7,0.9\}$ typically having the highest worst case and average performance. However, decreasing $\alpha$ by too much (i.e. $\alpha=0.1$) will result in a significant drop in performance, as will increasing $\alpha$ by too much ($\alpha=1$). This highlights the need for a careful tradeoff between dataset quality and diversity.
(a) Alpaca 3K
(b) Mixed 10K
(c) Ultrachat 10K
(d) LMSYS 10K
Figure 3: Effect of $\alpha$ on best case and worst case performance. The red line represents a randomly selected dataset. Worst reward refers to the bottom 10 percent of HH scores.
@@
Benchmark Performance. For a more comprehensive evaluation of QDIT, we follow InstructEval (Chia et al., 2023) and evaluate our model on various benchmark datasets including MMLU (Hendrycks et al., 2020), BBH (Suzgun et al., 2022), DROP (Dua et al., 2019), and HumanEval (Chen et al., 2021). Although these benchmarks are not fully aligned with instruction following performance, we find that QDIT uniformly improves MMLU performance and performs best on BBH and DROP in three out of the four scenarios when compared with quality driven selection. These results further confirm the efficacy of QDIT.
Table 2: Evaluation on benchmark datasets. We bold the best result out of quality based selection and QDIT based selection. Details can be found in Appendix G.
Alpaca 9K Mixed 10K
MMLU BBH DROP HumanEval MMLU BBH DROP HumanEval
Random 34.18 31.69 26.84 8.54 32.93 30.92 20.33 10.37
[…]
QDIT 35.64 30.85 26.55 7.93 33.29 32.36 26.19 10.37
Ultrachat 10K LMSYS 10K
MMLU BBH DROP HumanEval MMLU BBH DROP HumanEval
Random 32.12 33.19 26.24 13.41 33.05 32.57 25.06 15.85
[…]
QDIT 35.78 32.66 26.73 13.41 35.9 32.76 26.22 11.58
Different Base Models. We evaluate the performance of QDIT with different base models in Figure 3(a). From Figure 3(a) we find that QDIT can improve both average performance and worst case performance for other base models, supporting our main results. Interestingly we find that the Mistral model performs best according to HH score, but worst according to AlpacaFarm evaluation (Jiang et al., 2023). This highlights the importance of using multiple evaluation strategies.
(a) Different Base Models
(b) Different Data Sizes (Thousands)
@@
Figure 4: (a) Ablation with different base models. The train set is 10K instructions selected from Ultrachat. (b) Ablation with different data sizes. The worst reward is computed as the bottom 10 percent of HH scores, and the winning score is vs. Alpaca 52K.
@@
Data Size. We evaluate the performance of QDIT with different training data sizes in Figure 3(b), where we find that QDIT results in a 3-4 times improvement in data efficiency (and therefore training cost) as measured by robust performance. We also find that increasing the dataset size beyond $50K$ only results in marginal gains of average performance, and can even decrease worst case HH score. We hypothesize this phenomena is due to overfitting.
Case Study. We conduct a case study in Figure 5. We observe that quality based selection does not select an instruction similar to the test instruction and fails to respond correctly. On the other hand, QDIT does train on a similar instruction due to its more diverse training set and correspondingly responds to the instruction correctly. We remark that in practice the distance to the closest training instruction does not always correlate to evaluation performance, as we find embedding distance may not capture important aspects of an instruction such as the format it is in or the desired response format. In the future, we hope to investigate how different similarity metrics impact QDIT and instruction tuning performance.
Figure 5: Spotify playlist case study.
@@ 5 Conclusion
In this paper we introduce QDIT, a method for automatically selecting high quality and diverse instruction tuning datasets. QDIT allows us to flexibly tradeoff quality and diversity, improving both average performance and model robustness on a wide variety of datasets. As future lines of research, we would like to investigate how different similarity metrics affect the performance of QDIT and how QDIT performs with different model sizes.
Appendix A Tradeoff between Quality and Diversity: Additional Results
We display additional results on the effect of $\alpha$ on dataset diversity and quality in Figures 6, 7, 8. We again find that there is a tradeoff between quality and diversity, and that $\alpha$ can be used to control this tradeoff.
@@ Appendix A Tradeoff between Quality and Diversity: Additional Results
Appendix B Training Details
@@
Table 4: Selected values of $\alpha$.
Dataset Size Alpaca 52K Mixed 270K Ultrachat 1.3M LMSYS 1M
3K 0.7 - - -
9K 0.7 - - -
10K - 0.7 0.7 0.7
50K - 0.9 0.9 0.7
Appendix C QDIT Variants
In this section we describe and analyze the variants of QDIT.
Analysis. To further analyze why QDIT typically outperforms the two QDIT variants, we plot the selection strategy of each variant on a sample problem in Figure 9. From this figure we can observe the QDIT:cluster will sample equally from across the sample space, leading to some poor quality datapoints being sampled. We also observe that QDIT:Threshold does not necessarily optimize for diversity and may not sample from large portions of the sample space. On the other hand QDIT samples only high quality datapoints from each portion of the sample space, therefore resulting in a high quality and diverse dataset.
(a) QDIT:Cluster
(b) QDIT:Threshold
(c) QDIT
Figure 9: We display the selection strategy of the various QDIT algorithms on a test dataset. Six datapoints are selected and the selected datapoints are circled.
Appendix D Computational Cost of QDIT
@@ Appendix E Complete Win-Tie-Lose Results
In this appendix we show the win, ties, and losses achieved versus the reference models as judged by Claude 2. The results can be seen in Figures 10, 11, 12, 13, 14, and 15.
@@ Appendix G Benchmark Environments
For evaluation on common NLP benchmarks, we follow Chia et al. [2023]. In particular, we conduct five shot evaluation on MMLU, three shot on BBH, three shot on DROP, and zero shot on HumanEval.

### v3

@@ 4.3 Main Results
QDIT 36.13 32.12 60.71 26.73 69.8 86.8 52.05 37.34 32.52 61.44 26.41 69.28 85.0 52.0
Alpaca 3K Mixed 10K
MMLU BBH ARC DROP LAMBADA SCIQ AVG MMLU BBH ARC DROP LAMBADA SCIQ AVG
Random 36.17 30.25 61.67 26.32 71.64 87.0 52.18 32.93 30.92 58.34 20.33 68.1 84.1 49.12
@@
QDIT 35.47 30.44 61.95 27.02 69.68 84.1 51.44 34.29 31.23 60.71 26.00 69.72 89.8 51.96
Dolly 1K
MMLU BBH ARC DROP LAMBADA SCIQ AVG
Random 28.11 27.27 59.39 17.26 71.74 80.7 47.41
[…]
QDIT 33.78 30.33 59.84 22.59 72.26 80.6 49.9
@@ 4.4 Analysis
Effect of $\alpha$. We plot the effect of $\alpha$ on average and worst case performance in Figure 4. From Figure 4, we can see that the performance of QDIT changes smoothly with respect to $\alpha$, indicating that QDIT is relatively robust to the value of $\alpha$. In particular, values of $\alpha\in\{0.5,0.7,0.9\}$ typically having the highest worst case and average performance. However, decreasing $\alpha$ by too much (i.e. $\alpha=0.1$) will result in a significant drop in performance, as will increasing $\alpha$ by too much ($\alpha=1$). This highlights the need for a careful tradeoff between dataset quality and diversity.
@@
Benchmark Performance. For a more comprehensive evaluation of QDIT, we follow Chia et al. (2023) and Gao et al. (2021) by evaluating our model on various benchmark datasets including MMLU (Hendrycks et al., 2020), BBH (Suzgun et al., 2022), DROP (Dua et al., 2019), ARC (Clark et al., 2018), LAMBADA (Paperno et al., 2016), and SCIQ (Welbl et al., 2017). The results on these benchmarks can be found in Table 3. Details on our evaluation strategy can be found in Appendix G. Although these benchmarks are not fully aligned with instruction following ability, we find that QDIT typically improves benchmark performance compared to quality-driven selection, achieving a higher average score on four out of five datasets.
Different Base Models. We evaluate the performance of QDIT with different base models in Figure 6. From Figure 6 we find that QDIT can improve both average performance and worst case performance for other base models, including the more powerful LLama-2-13B. We remark that LLama-2-13B is only trained for 2 epochs, possibly resulting in a lower HH score.
(a) QDIT:Cluster
(b) QDIT:Threshold
(c) QDIT
Figure 5: The selection strategy of the various QDIT algorithms on an example dataset. Six data points are selected and the selected data points are circled.
@@
Figure 6: (a) Performance of QDIT with different base models. The dataset is Ultrachat. (b) Ablation with different data sizes. Worst HH Score refers to bottom 10% HH Score. The dataset is Ultrachat.
@@
Data Size. We evaluate the performance of QDIT with different training data sizes in Figure 6, where we find that QDIT leads to the highest gains in low-data regimes, but can still improve performance for larger datasets. We also find that increasing the dataset size beyond $50K$ only results in marginal gains of average performance, and can even decrease worst case HH score. This is likely due to the fact that the average reward score of the dataset decreases as the dataset size increases.
Figure 7: Case study on instruction generalization. This first column shows the test instruction, and the second column shows the most similar instruction in the respective training dataset. The third row shows the model outputs, where quality-based selection simply repeats the video requirements and QDIT-based selection provides useful video suggestions.
QDIT Variants. We study two variants of the QDIT algorithm that also attempt to improve data quality and diversity. The first variant, QDIT:Threshold, first orders the data point by quality, and then iteratively selects instructions that do not have a similarity score greater than $\tau$ with any point included selected subset. This algorithm essentially de-duplicates the dataset, and is similar to algorithms used in Wang et al. (2022b) and Liu et al. (2023). The second variant we propose is QDIT:Cluster, in which we first cluster the instructions and then select an equal amount of data points from each cluster in a quality-driven manner. Comprehensive details can be found in Appendix H.
We demonstrate how each variant selects algorithms in Figure 5. From these figures, we can see that the clustering-based approach may not work well as some clusters may be low quality, resulting in low-quality points being chosen. Moreover, the success of this variant is highly dependent on the success of clustering, which is difficult on large and imbalanced instruction tuning datasets. On the other hand, QDIT:Threshold will succeed in de-duplicating the dataset, but may fail to select a sufficiently diverse subset.
Experimentally, we observe in Figure 8(a) that the two variants of QDIT achieve slightly worse average performance compared to QDIT, but they often result in much worse worst-case performance.
(a) Performance of different QDIT variants on Ultrachat.
(b) Performance with different evaluators. The dataset is Alpaca.
Different Evaluators. To further investigate our finding that data diversity can improve instruction following performance, we compute the worst-case winning score with different evaluators. Concretely, we use Claude-2, Claude-2.1, GPT-3.5-Turbo, and GPT-4-Turbo as different model based evaluators (Achiam et al., 2023). We also conduct a blind human preference study, using the authors as annotators. Details on the human study can be found in Appendix F. The results using different models as evaluators can be seen in Figure 8(b). From these results we can see QDIT consistently improves worst-case performance, further validating our finding that data diversity improves robust instruction following capabilities.
Reducing the Train-Test Gap. One explanation for QDIT’s improvement in instruction following capabilities is a reduction in the gap between the training and testing data. More concretely, a model trained on a diverse dataset will be exposed to training instructions similar to those in the test set, and will therefore perform better on the test set.
A qualitative example of this phenomena can be found in Figure 7. In this example, the test instruction asks for help creating a short video. The closest example (measured by cosine similarity of the instruction embedding) in the quality-based dataset is unrelated to this task, while the closest instruction in the QDIT-based dataset is a similar instruction asking how to create a short video on nutrition. As an end result, the QDIT model provides useful video suggestions, while the Quality-driven model merely repeats the provided video requirements.
In all of our experiments we observe that QDIT selects datasets that better cover the testing dataset compared to quality-based selection. This can be seen in Figure 9, where we observe that the QDIT selects more similar instruction to the test instructions than quality-based selection does. This phenomenon provides one explanation on how QDIT improves instruction following ability, but in general it is difficult to directly attribute instruction following capabilities to test-set coverage.
Figure 9: Number of test instructions where quality based selection or QDIT contains the closest training example. Ties are not included.
@@ 5 Conclusion
In this paper we introduce QDIT, an automatic method for selecting high quality and diverse instruction tuning datasets. QDIT allows us to flexibly tradeoff quality and diversity, improving both average performance and model robustness on a wide variety of datasets. As a future line of research, we would like to investigate how different similarity metrics affect QDIT’s performance.
Appendix A Similarity Metrics
Appendix B Computational Cost of QDIT
Appendix C Tradeoff between Quality and Diversity: Additional Results
We display additional results on the effect of $\alpha$ on different datasets diversity and quality in Figures 10, 11, 12. We again find that there is a tradeoff between quality and diversity, and that $\alpha$ can be used to control this tradeoff.
@@ Appendix C Tradeoff between Quality and Diversity: Additional Results
Appendix D Training Details
@@
Table 5: The hyperparameter $\alpha$ used in the main experiments.
Data Size Dolly 15K Alpaca 52K Mixed 270K Ultrachat 1.3M LMSYS 1M
1K 0.7 - - - -
3K - 0.7 - - -
10K - - 0.9 - -
10K - - - 0.7 -
10K - - - - 0.7
@@ Appendix E Complete Win-Tie-Lose Results
In this appendix we show the win, ties, and losses achieved versus the reference models as judged by Claude 2. The results can be seen in Figures 13, 14, 15, 16, 17, and 18.
Appendix F Human Study Details
For the human study, we use a subset of the paper authors as annotators. For each example, we randomize the order of the reference model generation and evaluated model generation, to keep the study blind. We then ask the annotators to select their preferred generation according to the Alpacafarm prompt [Dubois et al., 2023].
@@ Appendix G Benchmark Environments
For evaluation on common NLP benchmarks, we follow Chia et al. [2023]. In particular, we conduct five shot evaluation on MMLU, three shot on BBH, and three shot on DROP. For ARC, LAMBADA, and SCIQ, we use the default zero shot setting of [Gao et al., 2021].
Appendix H QDIT Variants
In this section we describe and analyze the variants of QDIT.